# Pathwise uniform convergence of a full discretization for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise

### v4

@@ 5 Conclusions
In this paper, a novel stability estimate for a discrete stochastic convolution has been derived as an independent contribution. This stability estimate plays a crucial role in establishing pathwise uniform convergence estimates for fully discrete approximations of nonlinear stochastic parabolic equations. By combining this stability estimate with the discrete stochastic maximal $L^{p}$ regularity estimate, a pathwise uniform convergence rate of $O\big{(}h^{2-\epsilon}+\tau^{1/2}\big{)}$ has been achieved for a full discretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. Moreover, the convergence estimate employs general spatial $L^{q}$-norms. This theoretical finding not only advances our understanding of numerical methods for the stochastic Allen-Cahn equation but also introduces a novel approach for the numerical analysis of nonlinear stochastic parabolic equations, using the discrete stochastic maximal $L^{p}$ regularity estimate.
Funding
The first author was partially supported by Natural Science Foundation of Sichuan Province under grant 2023NSFSC1324. The second author was partially supported by National Natural Science Foundation of China under grant 12301525.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Declaration
Conflict of interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v5

@@ 5 Numerical results
In our theoretical framework, it is essential that the spatial domain $\mathcal{O}$ is a bounded, convex subset of $\mathbb{R}^{3}$ with a boundary of class $\mathcal{C}^{2}$. For the numerical experiments conducted in this study, we have chosen a cubic domain due to its simplicity and practicality, which facilitates both computational implementation and the interpretation of results. It is important to note that the theory of discrete stochastic maximal $L^{p}$-regularity remains an area of ongoing research.
@@ 6 Conclusions
In this paper, we have utilized discrete deterministic and stochastic maximal $L^{p}$ regularity estimates to derive a pathwise uniform convergence rate of $O\big{(}h^{2-\epsilon}+\tau^{1/2}\big{)}$ for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. This convergence estimate employs general spatial $L^{q}$-norms and imposes mild conditions on the diffusion coefficient. Our work introduces a new approach to establishing pathwise uniform convergence within the context of general spatial $L^{q}$-norms for finite element-based full approximations of nonlinear stochastic parabolic equations.